OpenStudy (anonymous):

simplify each expression: (3nsquared)cubed

OpenStudy (anonymous):



OpenStudy (mertsj):

Cube the two factors inside the parentheses. What is 3 cubed?

OpenStudy (anonymous):

27

OpenStudy (mertsj):

And what is (nsquared) cubed?

OpenStudy (anonymous):

i have no idea honestly. i dont want to incorrectly guess.

OpenStudy (anonymous):

i would say n to the 6 but , that sounds absurd .

OpenStudy (mertsj):

\[(n^2)^3 = n^2 \times n^2 \times n^2\]

OpenStudy (mertsj):

That is correct. n^6. Remember, when raising to a power, you multiply the exponents. So now you are done. 27n^6